Blade not working in latest build

Hi all,

I hope I put this in the correct place. The latest blade I assembled is not working in my current saber. I know that the blade is functional because it works in an older saber and I know that the current saber is functioning correctly because my original neopixel blade works in the current saber.

Some details:

  • current saber: Proffie V2.2 OS 7.8, Shtok V3 connector (short pins) wired in V4 (independent blade and connector)
  • current blade: SK6812 3535 strips with Shtok eco blade side pcb (no resistor on pad)
  • older saber: Proffie V1.5 OS 5.7, TCSS 7 pin connector
  • older blade: WS2812B 5050 strips with TCSS blade side pcb (330 ohm resistor on pad)

Neither saber has an additional resistor on the wire from the data 1 pad.

I thought that my connector alignment was the issue, but I was able to physically see the pins make contact with the data pad of the new blade when I removed the chassis from the hilt.
graflex2RGBv2_config.h (19.5 KB)
graflex2RGBv2 is the config for my older saber.
SSOBconfigV3.h (69.8 KB)
SSOBconfigV3 is for the current saber.

So what happens?
Does it just not light up at all?
Does it still make sounds, or does it stop working completely?

Yes that’s correct. Saber sounds and ignition/retraction are fine the blade just doesn’t light up. The Shtok blade connector lights up just fine.

Some Shtokd connectors comes with two resistors, and you’re supposed to remove one of them if you want to have one data wire go directly to the blade. If the resistor isn’t removed, you get two signals going to the blade, which can cause weird problems.

Not sure if that is the problem here, but it would probably be the first thing i would check.

Is it possible for you to access the wiring in the blade while connected to measure if data and voltage is present as it should be?

I was able to see from old pictures that I had removed the resistor from the R1 pad on the connector.

Unfortunately, I put hot glue over the solder joints in the blade connector and the exposed leads on the neopixel strips are covered in a few layers of heat shrink :sweat_smile:

I was able to check the resistance between the Data 1 and negative pads on the Proffie 1.5 and 2.2 and they both show ~450 ohms. Do you think the resistor on my older blade is making a difference? It’s strange how the new blade (no resistor) lights up just fine with the 1.5 setup.

Not sure why you would measure between those two places. I guess it’s good that they show the same value, but there shouldn’t be any resistor between those two.

Ultimately, I don’t really have any good clues as to what is going on here. I think it must be some sort of wiring issue, but it’s really hard to imagine something that would cause this particular symptoms. I suppose it could also be some incompatibility in the signal itself, but that’s equally weird to imagine.

In general, it’s either a data problem, a power problem, or gremlins.
I wrote some pages to try to help show how to test the first two problems:

Not sure what else I can do to help?